# MaxInfo: A Training-Free Key-Frame Selection Method Using Maximum Volume for Enhanced Video Understanding

### v2

@@ 4.2 Main Results
As Table 5 shows, the slow version outperforms the fast version in most cases, particularly when processing longer videos. The initial oversampling mechanism in the slow version provides MaxInfo with a richer selection space, which significantly improves performance. However, experimental results indicate that the fast version may outperform the slow version in specific cases, depending on the benchmark type and the number of initial samples. Nevertheless, in scenarios with limited computational resources or strict latency requirements, the fast version remains a practical and reliable choice due to its efficiency. This flexibility allows the two versions to adapt to different task requirements, offering users more options.
Table 5: Comparison of the impact of the slow and fast versions of the MaxInfo module on the performance of the LLaVA-Video-7B model on the LongVideoBench. All results are reproduced, except those marked with *, which are copied from the corresponding papers (Zhang et al., 2024c).
Model Input max. frames Output frames range Avg. frames Acc.
InternVL2 1B (Chen et al., 2024) 32 - 32 43.38
+ MaxInfofast 32 (1, 32) 30 43.6 +0.22%
+ MaxInfoslow 64 (1, 16) 16 44.65 +1.05%
InternVL2 2B (Chen et al., 2024) 32 - 32 46.97
+ MaxInfofast 32 (1, 32) 30 47.27 +0.73%
+ MaxInfoslow 64 (1, 16) 16 46.82 -0.15%
LLaVA-Video 7B (Zhang et al., 2024c) 64 - 64 58.2
+ MaxInfofast 64 (1, 64) 52 60.21 +2.01%
+ MaxInfoslow 128 (1, 64) 58 61.48 +3.28%
LLaVA-Video 72B (Zhang et al., 2024c) 64 - 64 61.9
+ MaxInfofast 64 (1, 64) 52 65.37 +3.47%
+ MaxInfoslow 128 (1, 64) 58 64.92 +3.02%
@@ 4.5.1 Vision Encoder Impact
Table 6 presents an ablation study on different visual encoders used for frame representation in MaxInfo.
Table 6: An ablation study on the choice of a visual encoder to obtain the visual representation of the frames in the MaxInfo approach. The best results are highlighed in bold, second best are underlined. Based on LLaVA-Video-7B (Zhang et al., 2024c)
Model Size Visual encoder Param. Acc.
LLava-Video 7B CLIP-ViT-Large 427.9 M. 58.94
LLava-Video 7B CLIP-ViT-Base 149.6 M. 58.79
LLava-Video 7B DINO2-base 86.6 M. 58.94
LLava-Video 7B DINO2-large 304.4 M. 58.86
LLava-Video 7B SigLIP-base-224 203.2 M. 59.76
LLava-Video 7B SigLIP-base-384 878 M. 59.24
We evaluated CLIP, DINOv2, and SigLIP as visual encoders. DINOv2 performed comparably to CLIP, despite lacking vision-language alignment. However, SigLIP outperformed both CLIP and DINOv2, likely due to its stronger language-vision connectivity. Interestingly, the larger SigLIP model underperformed, suggesting that more complex encoders require careful hyperparameter tuning (e.g., tolerance and rank) for optimal frame selection.
@@ 4.5.2 Impact of MaxInfo Block Hyperparameters
Effect of Different Tolerances and Ranks on the Model with Fixed Sampling. We conducted a series of tests with fixed benchmarks, model settings, and an initial pool of frames (e.g., $n^{*}=96$). As shown in Figure 6, the best observed result achieved a 3.3% improvement over the base LLaVA-Video-7B model with $\text{Tol}=0.3$ and $\text{R}=8$.
Figure 6: Effect of initial sampling on MaxInfo performance for LlaVa-Video 7B model (fixed R=8, Tol=0.15).
From these experiments, we derived the following guidelines:
@@
    Large Rank, Moderate Tol: When the Rank $R$ is large, a moderate Tolerance (Tol) helps avoid selecting overly similar frames, optimizing the trade-off between diversity and redundancy.
@@
    Balanced Selection: An excessively large $R$ combined with a very small Tol may cause the frame set to grow excessively, potentially exceeding the model’s capacity. Thus, $R$ and Tol should be selected based on the maximum feasible input frame count for a given LLM.
  3. 3.
    Performance Sensitivity: Our experiments show that performance is most sensitive to Tol values in the range $\text{Tol}\in[0.1,0.6]$. Beyond this range, improvements plateau or regress due to over-pruning or under-pruning.
  4. 4.
    Convergence: When setting $R=12$ and varying Tol values to 0.30, 0.45, and 0.60, the best result converged at 61.1%.
Effect of Sampling on Accuracy (Fixed Rank and Tolerance). We analyzed how varying the initial number of sampled frames impacts accuracy, while keeping the hyperparameters fixed (R=8, Tol=0.15). As shown in Figure 7, accuracy initially improves with the addition of more frames, but beyond a certain threshold, it begins to plateau or slightly decline. Our key observations are as follows:
Figure 7: Comparing the effect of initial sampling on the MaxInfo Block, n represents the initial number of frames sampled, and after passing through the MaxInfo Block, the number of frames is increased to a maximum of 64.
@@
    Increasing the initial number of frames provides more diverse information, enabling MaxVol to select better keyframes.
@@
    Beyond a certain number of frames, accuracy slightly declines due to excessive redundancy, which hampers MaxVol’s efficiency in keyframe selection.
  3. 3.
@@ 5 Conclusion
In this work, we introduced MaxInfo – a training-free method for selecting the most informative frames from videos, improving VLLMs inference. Our results consistently demonstrate that informative frame selection outperforms uniform sampling, leading to improved performance of state of the art VLLMs (LLaVA-Video, InternVL and Qwen2-VL with different sizes) across multiple benchmarks. However, the improvements are relatively small, which is expected given the training-free nature and the simplicity of our approach.
@@
Impact Statement
@@
This work introduces a training-free framework for improving frame sampling in Vision-Language Large Models (VLLMs), enhancing video understanding tasks. Such advancements are crucial for applications in education, accessibility, and public safety, where accurate and efficient video analysis can have a meaningful impact.
@@
However, improved video analysis capabilities may also raise ethical concerns, including potential misuse in surveillance, privacy violations, or biases in automated systems. Ensuring responsible deployment with a focus on fairness and transparency is essential to mitigate these risks.
@@
In summary, while our approach offers significant benefits, its adoption must adhere to ethical principles to promote equitable and responsible use.
Appendix A Appendix
A.1 Visualisation of MaxInfo Block results
Figure 8: MaxInfo Block sampled frames.
As shown Figure 8, MaxInfo Block can help us to reduce the redundancy of video information, better understand the video content and improve the model performance by selecting the most informative and representative key frames among multiple frames.
A.2 Additional comparisons with other models
We compared the performance of VLLMs of the same 7B size, as shown in Table 7, such as Open Source Model, Proprietary Models and Train-Free Model or Key Frames selection, on Video-MME Benchmark.
Table 7: Comparison of model results on Video-MME Benchmark.
Model LLM Size Frames Overall (WO/W-Subs) Short (WO/W-Subs) Medium (WO/W-Subs) Long (WO/W-Subs)
Proprietary Models
GPT-4o (OpenAI, 2024) - $384^{2*}$ 71.2/77.8 80.1/82.2 71.3/76.8 62.2/74.3
Gemini 1.5 Pro (Team et al., 2024) - 1/0.5 fps 75.0/81.3 81.7/84.5 74.3/81.0 67.4/77.4
Open Source Model
Video-LLaVA (Lin et al., 2023) 7B 8 39.9/41.6 45.3/46.1 38.0/40.7 36.2/38.1
Long-LLaVA (Wang et al., 2024b) 7B 64 52.9/57.1 61.9/66.2 51.4/54.7 45.4/50.3
mPLUG-Owl3 (Ye et al., 2024) 7B 128 59.3/68.1 70.0/72.8 57.7/66.9 50.1/64.5
LongVA (Zhang et al., 2024b) 7B 128 52.6/54.3 61.1/61.1 50.4/53.6 46.2/47.6
LongVILA (Xue et al., 2024) 7B 128 57.5/60.6 69.3/70.4 56.1/59.2 47.0/52.1
256 57.4/61.8 69.0/70.8 56.7/60.6 46.4/54.0
Train-Free Model or Key Frames selection
Video-Tree* (Wang et al., 2024c) - - -/- -/- -/- 54.2/-
LongVU* (Shen et al., 2024) 7B 1fps -/- -/- -/- -/60.6
Chat-UniVi (Jin et al., 2024a) 7B 64 40.6/- 45.7/- 40.3/- 35.8/-
MaxInfo Method
Qwen2-VL-MaxInfo 7B Avg.180 62.1/70.0 72.5/76.4 62.0/69.8 51.8/63.9
LLaVA-Video-MaxInfo 7B Avg.58 64.2/71.4 74.6/78.7 63.3/69.8 54.6/65.9
A.3 Impact of MaxInfo Block’s Tol and Rank on Models.
Overview of the details of how Tol and R affect the accuracy of the model, as Table 8, I.Frame is the number of frames input to the MaxInfo module, and F.Frame is the maximum number of frames after going through the MaxInfo Block. The effect of the initial number of samples on the model as shown in Table 9.
Table 8: Comparison of experimental results on the effect of different tol and rank on the LLava-Video (Zhang et al., 2024c).
Model Size R Tol (I.Frames, F.Frames) Acc.
LLava-Video 7B 12 0.15 (96, 64) 60.35
LLava-Video 7B 12 0.30 (96, 64) 61.25
LLava-Video 7B 12 0.45 (96, 64) 61.10
LLava-Video 7B 12 0.60 (96, 64) 61.14
LLava-Video 7B 8 0.15 (96, 64) 60.06
LLava-Video 7B 8 0.30 (96, 64) 61.48
LLava-Video 7B 8 0.45 (96, 64) 59.68
LLava-Video 7B 8 0.60 (96, 64) 57.80
LLava-Video 7B 6 0.10 (96, 64) 61.10
LLava-Video 7B 6 0.15 (96, 64) 60.96
LLava-Video 7B 6 0.30 (96, 64) 60.58
LLava-Video 7B 6 0.45 (96, 64) 59.01
LLava-Video 7B 6 0.60 (96, 64) 57.89
Table 9: Comparison of the effect of different initial sampling values on the model.
Model Size (I.Frames, F.Frames) Acc.
LLava-Video 7B (256, 64) 60.58
LLava-Video 7B (192, 64) 60.73
LLava-Video 7B (128, 64) 60.88
LLava-Video 7B (96, 64) 60.06

### v3

@@ 4.1 Main Results
As shown in Table 4, the slow version outperforms the fast version in most cases, especially when processing long videos. Its initial oversampling mechanism provides MaxInfo with a richer selection space, which significantly improves performance. However, experiments have also shown that the fast version may outperform the slow version in certain benchmarks or when the number of initial samples is small. The MaxInfo block time is very short, so latency is almost negligible and more details on latency and GPUs memory and time.
Table 4: Performance of Fast vs. Slow MaxInfo variants on LLaVA-Video-7B (LongVideoBench). I/O: input/output frames. All results are reproduced, except those marked with *, which are copied from the corresponding papers [llava-video].
Model I max. frames O frame range Avg. frames Encoding + MaxInfo Time Acc.
InternVL2 1B [internvl] 32 - 32 - 43.38
+ MaxInfofast 32 (1, 32) 30 0.179 + 0.0126 s 43.60 +0.22%
+ MaxInfoslow 64 (1, 16) 16 0.339 + 0.0215 s 44.65 +1.05%
InternVL2 2B [internvl] 32 - 32 - 46.97
+ MaxInfofast 32 (1, 32) 30 0.179 + 0.0126 s 47.27 +0.30%
+ MaxInfoslow 64 (1, 16) 16 0.339 + 0.0215 s 46.82 -0.15%
LLaVA-Video 7B [llava-video] 64 - 64 - 58.2
+ MaxInfofast 64 (1, 64) 52 0.339 + 0.0215 s 60.21 +2.01%
+ MaxInfoslow 128 (1, 64) 58 0.624 + 0.0421 s 61.48 +3.28%
LLaVA-Video 72B [llava-video] 64 - 64 - 61.9
+ MaxInfofast 64 (1, 64) 52 0.339 + 0.0215 s 65.37 +3.47%
+ MaxInfoslow 128 (1, 64) 58 0.624 + 0.0421 s 64.92 +3.02%
@@ 4.3.1 Vision Encoder Impact
As shown in Table 5, we evaluated CLIP, DINOv2, and SigLIP as visual encoders. DINOv2 performed comparably to CLIP, despite lacking vision-language alignment. However, SigLIP outperformed both CLIP and DINOv2, likely due to its stronger language-vision connectivity. Interestingly, the larger SigLIP model underperformed, suggesting that more complex encoders require careful hyperparameter tuning (e.g., tolerance and rank) for optimal frame selection.
Table 5: Visual encoder ablation for MaxInfo. Best in bold, second-best underlined. Based on LLaVA-Video-7B [llava-video].
Model Visual Encoder Param. Acc.
LLaVA-Video [llava-video] CLIP-ViT-Large 427.9M 58.94
LLaVA-Video [llava-video] CLIP-ViT-Base 149.6M 58.79
LLaVA-Video [llava-video] DINO2-base 86.6M 58.94
LLaVA-Video [llava-video] DINO2-large 304.4M 58.86
LLaVA-Video [llava-video] SigLIP-base-224 203.2M 59.76
LLaVA-Video [llava-video] SigLIP-base-384 878.0M 59.24
@@ 4.3.2 Impact of MaxInfo Block Hyperparameters
Effect of Different Tolerances and Ranks on the Model with Fixed Sampling. We conducted a series of tests with fixed benchmarks, model settings, and an initial pool of frames (e.g., $n^{*}=96$). As shown in Figure 3, the best observed result achieved a 3.3% improvement over the base LLaVA-Video-7B model with $\text{Tol}=0.3$ and $\text{R}=8$. From these experiments, we derived the following guidelines:
@@
    Performance Sensitivity: Our experiments show that performance is most sensitive to Tol values in the range $\text{Tol}\in[0.15,0.60]$. Beyond this range, improvements plateau or regress due to over-pruning or under-pruning.
@@
    Convergence: The model will converge when setting $\text{R}\in[12,15],\text{Tol}\in[0.3,0.45]$, this means that our choice of hyperparameters is not intractable.
Figure 3: Effect of initial sampling on MaxInfo performance for LlaVa-Video 7B model.
Effect of Sampling on Accuracy We analyzed how varying the initial number of sampled frames impacts accuracy, while keeping the hyperparameters fixed (R=8, Tol=0.15). As shown in Figure 4, accuracy initially improves with the addition of more frames, but beyond a certain threshold, it begins to plateau or slightly decline. Our key observations are as follows:
@@
    Increasing the initial number of frames provides more diverse information, enabling MaxVol to select better keyframes, but the information converges.
@@
Figure 4: Effect of Initial Sampling on MaxInfo. Starting from $n^{*}$ sampled frames, the MaxInfo Block selects up to 64 informative frames for further processing.
4.3.3 Case Study
Figure 5 presents a long-video example from the Video-MME [video-mme] dataset, qualitatively illustrating the effectiveness of the proposed MaxInfo Block. We compare frame selection by MaxInfo Block with Uniform Sampling and observe that the keyframes chosen by MaxInfo are more closely aligned with the manually annotated Ground Truth–related frames.
Figure 5: Qualitative: (a) MaxInfo vs Uniform Sampling with GT-aligned frames; (b) CLIP scores show MaxInfo’s answer coverage in single samples.
@@ 5 Conclusion
In this work, we introduced MaxInfo – a training-free method for selecting the most informative frames from videos, improving VLLMs inference. Our results consistently demonstrate that informative frame selection outperforms uniform sampling, leading to improved performance of state of the art VLLMs (LLaVA-Video, InternVL and Qwen2-VL with different sizes) across multiple benchmarks. For example, MaxInfo achieves a 3.28% improvement on LongVideoBench and a 6.4% improvement on EgoSchema for LLaVA-Video-7B. The Slow Version of MaxInfo improves LLaVA-Video-72B performance by 3.47% on LongVideoBench.
@@
6 Acknowledgments
Innopolis University authors were supported by the Research Center of the Artificial Intelligence Institute at Innopolis University. Financial support was provided by the Ministry of Economic Development of the Russian Federation (No. 25-139-66879-1-0003).
Appendix A Implementation Details
We mostly focus on longer videos because better frame selection plays a bigger role in longer, more complex videos, whereas shorter ones intuitively work well with uniform sampling due to their lower information content and complexity.
For the general multiple-choice question-answering evaluation, we follow the official guidelines to construct the instructions using the provided questions and options. We added a prompt to the question and options like ”Respond with only the letter (A, B, C, or D) of the correct option.” for LongVideoBench [longvideobench], Video-MME [video-mme], MLVU [mlvu] and MVBench [mvbench] or ”Answer with the option’s letter from the given choices directly and only give the best option.” for EgoSchema [egoschema]. We follow the original benchmarks setup to calculate the final scores, and we also align our evaluation protocols with other evaluation toolkits, such as lmms-eval [zhang2024lmmseval].
Appendix B Additional Experiments and Details
To further assess the impact of MaxInfo, we evaluate its performance with an additional set of models [wang2025internvl3], [yao2024minicpmv] on the LongVideoBench and Video-MME benchmarks.
B.1 Applying MaxInfo to recent models
The results in Table 6 show that MaxInfo consistently improves model performance across both benchmarks, suggesting that precise frame selection is particularly important for long-video tasks.
Table 6: Adaptation of MaxInfo to current new long video understanding models
Model Size Frame Interval Avg. frames LongVideoBench.
MiniCPM [yao2024minicpmv] 9B 128 128 56.17
+ MaxInfo 9B [8, 82] 56 59.61
[…]
InternVL3.5 [wang2025internvl3] 1B 16 16 47.7
+ MaxInfo 1B [1, 16] 16 49.0
[…]
InternVL3.5 [wang2025internvl3] 8B 16 16 57.4
+ MaxInfo 8B [1, 16] 16 59.0
[…]
InternVL3.5 [wang2025internvl3] 38B 16 16 60
+ MaxInfo 38B [1, 16] 16 61.6
[…]
B.2 Performance Analysis: MaxInfo vs. Uniform Sampling
To better understand the strengths and trade-offs of MaxInfo, we analyzed per-task accuracy across multiple benchmarks. Our results, as shown in Figure 6, indicate that MaxInfo performs superiorly in high information density tasks such as counting, summarizing and spatial reasoning, while uniform sampling has a slight advantage in tasks that rely on temporal continuity, reflecting the key trade-off between information maximization and temporal consistency.
Figure 6: Accuracy comparison between Uniform Sampling and MaxInfo across three benchmarks.
B.3 Comparison with CLIP baseline
As shown in Table 7, we compare the experimental results of two keyframe extraction strategies based on the QwenVL2-2B model on the LongVideoBench benchmark: the CLIP-Based thresholding method and the MaxInfo module method. Both methods extract the same number of frames in the initial phase, so the encoding time is kept the same, where the similarity threshold of the CLIP-Based method is set to 0.5. The results show that the MaxInfo module outperforms the CLIP-Based method in terms of the overall performance in keyframe selection.
Table 7: Performance comparison on LVBench.
Model Method Accuracy
QwenVL2-2B CLIP-Based 44.3
QwenVL2-2B CLIP-Based + MaxInfo 44.5
QwenVL2-2B MaxInfo + CLIP-Based 43.8
QwenVL2-2B MaxInfo 48.8
In addition, we also explored combining the CLIP-Based method with MaxInfo module. The experiments show that MaxInfo is able to improve the overall information quality of the input sequences, and its information maximization strategy plays a key role in frame selection, which further enhances the performance of the model. CLIP-Based loses a lot of semantic information, which can lead to performance degradation of the model.
In order to further evaluate whether MaxInfo will lose the key frames related to the problem, we compare MaxInfo with the Uniform Sampling method under the CLIP Score metric. The experimental results shown in Table 8 that MaxInfo does not miss the frames related to the semantics of the problem, and is able to retain the semantic relevance effectively.
Table 8: CLIP score comparison between uniform and MaxInfo sampling.
Sampling Method CLIP-score
Uniform 0.37
MaxInfo 0.39
B.4 Qualitative comparison with uniform sampling
We randomly selected 50 video samples in LongVideoBench and calculated the cosine similarity between the frames selected by MaxInfo and the cosine similarity between the frames obtained by uniform sampling.
Figure 7 shows the distribution of cosine similarity for the same number of frames. It is clear that MaxInfo produces a more diverse distribution like a low similarity offset compared to uniform sampling, highlighting its ability to capture more diverse visual content.
Figure 7: Similarity distribution between neighbouring frames ($frame_{i}$ and $frame_{i+1}$).
As shown in Figure 8, we plotted 200 sampled data points to improve visual clarity. The results show that our MaxInfo module exhibits higher diversity in frame selection compared to uniform sampling.
Figure 8: CLIP similarity between neighboring frames selected by MaxInfo module.
B.5 Computational Efficiency: Time and Memory Consumption
When processing long videos, the LLM is the most resource-intensive component of VLLMs due to its parameter count and the quadratic complexity of attention with respect to input length. Since most of the context is occupied by visual tokens from frames, our MaxInfo method reduces this load by selecting keyframes. Importantly, MaxInfo requires minimal and constant memory and preprocessing time, independent of LLM size, and remains significantly lighter than uniformly sampling all frames.
Time Complexity. To evaluate the latency overhead of MaxInfo in practice, we measured its runtime with the Qwen2-VL model. As shown in Table 9, the runtime of MaxInfo is almost negligible compared to the inference time of the VLLM itself, confirming that MaxInfo is a lightweight and efficient frame selection mechanism. The initial VLLM time means the inference time of the 512 frames of information directly into the Qwen2-VL model. The frame count selected by MaxInfo Block is adaptive to the information content of the input. For near-static videos (low information density), MaxInfo drastically reduces the number of processed frames. Consequently, VLLMs + MaxInfo Block may achieve lower time compared to the initial VLLMs + MaxInfo Block configuration. The times reported in the table represent an upper bound; in practice, the reduced number of frames can lead to several-fold speedups on certain tasks. All experiments were conducted on an A100 GPU.
Table 9: Runtime of different pipeline components, based on Qwen2-VL. Frame size = 512 (UP is Upper Bound).
Model Size CLIP (s) MaxVol (s) VLLMs (s) VLLMs + MaxInfo (UP)
2B 0.296 0.0109 2.979 $\leq$ 3.285
7B 0.296 0.0109 5.372 $\leq$ 5.679
72B 0.296 0.0109 30.737 $\leq$ 31.044
We also analyzed the running time of the MaxVol algorithm alone, including its chunk-based variant, under different initial numbers of frames, as shown in Table 10. The experimental results show that the running time of MaxVol remains low across settings, with minimal impact on the overall inference efficiency.
Table 10: MaxVol algorithm runtime (excluding image encoding time) for different input sizes.
Method Input Size MaxVol Time (s)
MaxInfo 128 0.0044
MaxInfo 256 0.0053
MaxInfo 512 0.0109
Chunks-Based MaxInfo $32\times 32$ 0.0375
Then we estimated CUDA inference time across different VLLM sizes which is shown in Figure 9. The overhead of MaxInfo remains small and nearly constant, while the overall inference time grows with model size, demonstrating that MaxInfo adds minimal cost compared to the savings from reduced visual tokens. For small models (up to 8B parameters), the relative benefit is limited since inference cost is low. However, for larger models (26B–76B), MaxInfo provides clear efficiency gains by substantially reducing the number of visual tokens, making its impact especially pronounced for long-video tasks where input length dominates computational cost.
Figure 9: CUDA inference time across different VLLM sizes. The preprocessing cost of MaxInfo remains small and nearly constant, while overall inference time increases with model size.
Figure 10: The comparison for memory performance on the GPU for InternVL2 models with and without MaxInfo module. The dashed line shows the CUDA memory requirements for MaxInfo.
Memory Consumption. Secondly, we precisely evaluated memory consumption of out approach. As shown in Figure 10, MaxInfo’s CUDA memory usage remains constant for a fixed number of initial frames and grows much more slowly than uniform sampling as LLM size increases.
In summary, our analysis of time and memory efficiency shows that MaxInfo introduces only negligible overhead while substantially reducing the computational burden of processing long videos. Its constant preprocessing cost and slower growth in memory usage make MaxInfo particularly advantageous for large-scale VLLMs, with the benefits becoming most pronounced for models exceeding 10B parameters.
Appendix C Theoretical Justification
Definition 1.
Definition of the maximum volume of the video frame feature matrix.
We consider a matrix $Q\in\mathbb{C}^{N\times r}$, where each row represents the CLIP or SigLIP etc. feature of a video frame, ordered sequentially in time, where N denotes the number of frames and r denotes the dimension of the feature.
We aim to identify a submatrix $\hat{Q}\in\mathbb{C}^{K\times r}$ of the original matrix $S$, such that $\hat{S}$ closely approximates $S$ in terms of matrix volume, thereby preserving its essential structural information.
To obtain the submatrix $\hat{Q}$, we introduce a coefficient matrix $C$ based on the minimum-norm linear combination as Equation 10
$\tilde{C}\hat{Q}=\tilde{Q}$ (10)
Here, $\tilde{Q}$ denotes a set of sample rows selected from the original matrix $Q$ for reconstruction. By solving for $\tilde{C}$, we can approximate the reconstruction of $\tilde{Q}$ using only the representative rows in $\hat{Q}$. In addition, it is shown that the selected K rows are the most representative of the video frame information.
Solving. The submatrix $\hat{Q}\in\mathbb{C}^{K\times r}$ provides an approximation of the original matrix $Q\in\mathbb{C}^{N\times r}$ within a tolerance $\tau$.
We start with an initial submatrix $\hat{Q}\in\mathbb{C}^{M\times r}$ and add a row $Q_{i}\in\mathbb{C}^{1\times r}$ to each iteration to bring the expanded submatrix up to speed in the sense of volume. The updating process can be expressed as follows Equation 11 and the volume of the updated matrix can be defined as Equation 12
$\hat{Q}\leftarrow\begin{bmatrix}\hat{Q}\\ Q_{i}\end{bmatrix}$ (11)
$\text{Vol}(\hat{Q})_{\text{new}}=\text{Vol}(\hat{Q})_{\text{old}}\cdot\sqrt{1+\|\tilde{C}_{i}\|_{2}^{2}}$ (12)
where $Q_{i}$ is the row selected from the original matrix $Q\in\mathbb{C}^{N\times r}$ that currently boosts the volume of the submatrix the most. Repeat this process iteratively until the conditional Equation 13 is satisfied or the target number of $K$ rows is reached.
$\|\tilde{C}_{i}\|_{2}\leq\tau$ (13)
Proof of maximum information entropy. To justify our approach, we use differential entropy as an information measure. Suppose our normalized frame embeddings form a matrix $S$. The differential entropy of a uniform distribution over the convex hull $\mathcal{C}(S)$ is given by the following Equation 14.
$H_{\max}(S)=\ln(\text{Vol}(\mathcal{C}(S)))$ (14)
where $\text{Vol}(\mathcal{C}(S))$ is the volume of the convex hull formed by selected embeddings. Classical results show the following Equation 15.
$\text{Vol}(\mathcal{C}(S))=\kappa\sqrt{\det(S^{\top}S)}$ (15)
for some constant $\kappa>0$. Thus we get Equation 16
$H_{\max}(S)=\ln V(S)+\text{constant}$ (16)
where $V(S)=\sqrt{\det(S^{\top}S)}$. Since MaxVol maximizes $V(S)$, it maximizes the upper bound on differential entropy, ensuring that selected frames are more informative.
In summary, we can theoretically select the most representative frame information. The feature matrices corresponding to the selected frames have good linear independence under the constraint of the tolerance parameter $\tau$, thus constituting an approximately optimal subset of the representation. This process achieves our goal of information maximization, i.e. preserving the most critical structural information while compressing redundancy.
Appendix D Societal Impacts
@@
This work introduces a training-free framework for improving frame sampling in Vision-Language Large Models (VLLMs), enhancing video understanding tasks. Such advancements have important implications for applications in education, accessibility, and public safety.
@@
However, improved video analysis capabilities may also raise ethical concerns, including potential misuse in surveillance, privacy violations, or biases affecting different communities. Ensuring responsible deployment with fairness and transparency is essential to mitigate these risks.
@@
In summary, while our approach provides significant benefits, its adoption should adhere to ethical principles to promote equitable and responsible use.